OpenStudy (anonymous):

PLEASE HELP!! Find the amplitude y = -5 sin 1/2x.

jimthompson5910 (jim_thompson5910):

In general, the amplitude of y = A*sin(Bx - C) + D is |A|

jimthompson5910 (jim_thompson5910):

Ex: the amplitude of y = 8*sin(3x-2) is |8| = 8 the amplitude of y = -14*sin(pi*x)-4 is |-14| = 14

OpenStudy (anonymous):

Thank you so much!!!!! your a life saver!

jimthompson5910 (jim_thompson5910):

you're welcome